# Policy Learning with Confidence

### v1

@@ 3.2.1. Immunization Nudges
The outcomes include the number of measles shots administered, and number of shots per dollar spent. We observe 7,370 households and use doubly robust scores to obtain the welfare estimates $\widehat{V}(\pi)$ and their standard errors. Following Banerjee et al. (2024), we weight these village-level regressions by village population, and standard errors are clustered at the SC level. In light of sampling weights and clustered dependence, developing a suitable Gaussian bootstrap procedure can be complicated. For simplicity, we use the Bonferroni-corrected $\widehat{k}=\Phi^{-1}\bigl{(}1-\alpha/\lvert\mathcal{T}\rvert\bigr{)}$, where $\Phi(\cdot)$ denotes the standard normal CDF. This method still guarantees a valid high-confidence lower bound.
@@
Panel A. Outcome: Measles shots. Control:
@@
LCB: $\widehat{V}(\widehat{\pi})-\widehat{q}_{0.95,\Pi}\widehat{s}(\widehat{\pi})$ 6.56 7.8 2.05
@@
Panel C. Optimal Treatment Policy:
@@
Note. EWM refers to empirical welfare maximization, and PoLeCe corresponds to policy learning with confidence. In Panels A and B, the EWM and PoLeCe lower confidence bounds with $\alpha=0.05$ are provided. In Panel C, optimal treatment policies are given.
@@ 3.2.2. Informal Savings Technologies
Panel B. Optimal Treatment Policy:
@@
Note. EWM refers to empirical welfare maximization, and PoLeCe corresponds to policy learning with confidence. In Panel A, EWM and PoLeCe lower confidence bounds with $\alpha=0.05$ are provided along with that for the control group policy. In Panel B, an optimal treatment policy for each vector of (Female, Married) is given.
There are $|\mathcal{T}|=624=5^{4}$ possible treatment policies all together because there are five treatment values and four possible values in the support of the two covariates. Figure 3(a) plots all welfare estimates relative to their precision. The highest welfare is observed below 900 Ksh, but it is accompanied by the largest standard error. Therefore, it is advisable to account for the precision of these estimates when selecting an optimal policy. To implement PoLeCe at the level $\alpha=0.05$ and calculate $\widehat{q}_{0.95,\Pi}$, we apply the bootstrap procedure from (B). Table 6 summarizes the empirical results. In Panel A, the maximized welfare by EWM is 890 Ksh, which corresponds to the highest welfare in Figure 3(a). The adjusted estimate of maximum welfare $\widehat{V}(\widehat{\pi}_{\mathrm{PoLeCe}})$ at $\alpha=0.05$ is 587, which is much less than no-precision-corrected $\widehat{V}(\widehat{\pi}_{\text{EWM}})$ but substantially greater than $\widehat{V}(\widehat{\pi}_{\text{EWM}})-\widehat{q}_{0.95,\Pi}\widehat{s}(\widehat{\pi}_{\text{EWM}})$. Furthermore, the chosen optimal policies differ for females. EWM selects Health Pot for unmarried women and Safe Box for married women. In contrast, after accounting for standard errors, PoLeCe recommends Safe Box for unmarried women and Health Pot for married women.
@@ 3.2.3. Alcohol and Self-Control
Panel B. Optimal Treatment Policy:
@@
Note. EWM refers to empirical welfare maximization, and PoLeCe corresponds to policy learning with confidence. In Panel A, the EWM point estimate and the PoLeCe with $\alpha=0.05$ are provided along the control group policy. In Panel B, an optimal treatment policy for each vector of (Baseline sober, Owns rickshaw) is given.
There are $|\mathcal{T}|=81$ possible treatment policies in total. This follows from $81=3^{4}$, given three treatment levels and four possible values of the two covariates. Figure 3(b) plots the welfare estimates against their precision. Although the highest welfare appears with a relatively small standard error, this does not necessarily imply statistical significance. To implement PoLeCe at the level $\alpha=0.05$ and calculate $\widehat{q}_{0.95,\Pi}$, we apply the bootstrap procedure from (B). Table 7 summarizes the empirical results. The EWM and PoLeCe estimates, along with their lower confidence bounds, are very similar and only slightly exceed those of the control group.
@@ 4. Conclusion
In this paper we have focused on the problem faced by a DM who wishes to choose from a menu of policies that which will deliver the highest benefit using imperfect sample estimates. In order to balance the estimated performance of each policy with its associated statistical uncertainty, we proposed and analyzed the properties of a class of risk-aware policies that make the precision/performance tradeoff explicit. Such policies solve the expected loss function of a risk-averse planner and achieve favorable regret properties. We proposed a specific rule from the class of risk-aware policies, namely the PoLeCe rule, which uses a data-dependent construction for balancing the inherent tradeoff between estimated performance and sample uncertainty, such that a lower confidence bound on the welfare obtained by the chosen policy is automatically provided. Application of the rule was demonstrated across four settings; one featuring the allocation of government funds across social programs using the MVPF, and three others featuring optimal treatment assignment. Intuitively, the approach offers a principled and robust way for a decision maker to hedge against selecting the best-performing policy in sample, when that policy exhibits large sample variation relative to other well-performing policies. Importantly, risk-aware policy rules allow policies with fractional allocations.
@@
A large body of work synthesized in Manski (2013, 2019, 2024) has advocated for greater acknowledgement and incorporation of uncertainty in planning problems. This paper contributes by proposing a principled way to acknowledge and incorporate statistical uncertainty into decision making. Statistical uncertainty is however only one of the many different types of uncertainty that may be present. [^20] For example, Manski (2019) discusses transitory uncertainty, permanent uncertainty, and conceptual uncertainty. The statistical uncertainty considered here is one type of permanent uncertainty. We have focused on settings in which the benefit of each policy is point identified, such that consistent estimates of the benefit and sample variance of the policies are available. Application of the concepts developed here to settings that feature so-called “deep uncertainty”, or ambiguity, would seem an important direction for further development given the large number of settings in which the mean performance of various policies is only credibly partially identified. [^21] There is now a long line of research on treatment choice when mean treatment performance is only partially identified, going back to Manski (2000). Recent research with references to the broader literature includes Russell (2020), Ishihara and Kitagawa (2021), Yata (2021), Christensen et al. (2023), Kido (2023), and Montiel Olea et al. (2023). This would require balancing statistical uncertainty with a collection of interval estimates for each policy’s performance.
@@
While the nuance required to formally develop such an approach is beyond the scope of the present paper, a rough prescription could be made based on an extension of the reporting guarantee established here. To see how, consider the definition of the PoLeCe rule in (9). The rule selects the policy that maximizes a $1-\alpha$ lower confidence band for the maximal achievable welfare. Construction of such a confidence band does not require that the optimal level of welfare or the optimal rule be point identified, and could be implemented using techniques developed in e.g. Chernozhukov et al. (2013) under partial identification. This is of course not the only way to balance statistical uncertainty and ambiguity. A more thorough study of the performance of such an approach, and the implicit tradeoff between statistical uncertainty and ambiguity due to partial identification would seem a useful direction for future research.
@@ Appendix A Deferred Proofs
Proof of Proposition 1
Claim (1). Let $v=\widehat{V}(\pi)$ and $s=\widehat{s}(\pi)$. Then
[…]
This maximization can be written in a Lagrangian form:
Proof of Proposition 2.
Observing that
and
the program becomes
@@
and by the standard calculatiom we have $\mathrm{E}\sup_{\pi\in{K}}Z_{\pi}\leq\sqrt{2\log|K|}$, where $|{K}|\leq p$ is the cardinality of ${K}$. ∎
@@
$\bar{\sigma}_{\Pi}:=\sqrt{n}\max_{\pi\in\Pi}\widehat{s}(\pi);\quad\widehat{\pi}:=\widehat{\pi}_{RW}(\widehat{k});$
@@
$V_{\max}-V(\widehat{\pi})$$\leq$ $V_{\max}-L(\widehat{k})+L(\widehat{k})-V(\widehat{\pi})$
@@ Appendix B Efficient Decision Frontier Computation
For the investment allocation problem, it may be infeasible to discretize $\Pi$ and apply the above algorithm to plot the efficiency frontier. Instead, we start with a fine grid for $k$ centered at $\hat{q}_{0.95,\Pi},$ and directly trace out the efficiency decision frontier by solving the corresponding opitimization problem as in  (4). Appendix C provides computationally efficient way to solve such optimization and therefore to plot the efficiency frontier.
@@ Appendix C Gaussian Bootstrap for Investment Allocation Problem
$f(t)=\max_{\pi\in\Pi}\left(\pi^{\top}Z^{*}-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}\right).$
The following lemma formalizes the properties of $f(t)$ necessary for this approach.
@@
Let $\Pi\subset\mathbb{R}^{d}$ be a convex, bounded subset of the probability simplex, and let $\widehat{\Omega}$ be a positive-definite matrix. Define
@@
which is finite since $\Pi$ is bounded. As $t\to\infty$, the term $-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}$ dominates for every $\pi\in\Pi$, leading to $f(t)\to-\infty$. Since $f(t)$ is continuous (as the pointwise maximum of continuous functions over a compact set), the Intermediate Value Theorem guarantees the existence of some $t^{*}\geq 0$ such that $f(t^{*})=0$.
@@ Appendix D Further Details on the MVPF Estimates
For ease of reference, the Policy Impacts Library (Hendren and Sprung-Keyser (2021)) descriptions of the social programs corresponding to each MVPF estimate in Table 2 are provided below.

### v2

@@ 3.2.1. Immunization Nudges
The outcomes include the number of measles shots administered, and number of shots per dollar spent. We observe 7,370 households and use doubly robust scores to obtain the welfare estimates $\widehat{V}(\pi)$ and their standard errors. Following Banerjee et al. (2024), we weight these village-level regressions by village population, and standard errors are clustered at the SC level. We calculate $\widehat{k}=\widehat{q}_{1-\alpha,\Pi}$ by the bootstrap procedure in Condition (B).
@@
Panel A. Outcome: Measles shots
@@
LCB: $\widehat{V}(\widehat{\pi})-\widehat{q}_{0.95,\Pi}\widehat{s}(\widehat{\pi})$ 6.57 7.81 2.06
@@
Panel C. Selected Policy:
@@
Note. EWM refers to empirical welfare maximization, and PoLeCe corresponds to policy learning with confidence. In Panels A and B, the EWM and PoLeCe lower confidence bounds with $\alpha=0.05$ are provided. In Panel C, selected treatment combinations are given.
@@ 3.2.2. Informal Savings Technologies
Panel B. Selected Policy:
@@
Note. EWM refers to empirical welfare maximization, and PoLeCe corresponds to policy learning with confidence. In Panel A, EWM and PoLeCe lower confidence bounds with $\alpha=0.05$ are provided along with that for the control group policy. In Panel B, selected treatment for each vector of (Female, Married) is given.
There are $|\mathcal{T}|=625=5^{4}$ possible treatment policies all together because there are five treatment values and four possible values in the support of the two covariates. Figure 3(a) plots all welfare estimates relative to their precision. The highest welfare is observed below 900 Ksh, but it is accompanied by the largest standard error. Therefore, it is advisable to account for the precision of these estimates when selecting an optimal policy. To implement PoLeCe at the level $\alpha=0.05$ and calculate $\widehat{q}_{0.95,\Pi}$, we apply the bootstrap procedure from Condition (B). Table 5 summarizes the empirical results. In Panel A, the maximized welfare by EWM is 890 Ksh, which corresponds to the highest welfare in Figure 3(a). The adjusted estimate of maximum welfare $\widehat{V}(\widehat{\pi}_{\mathrm{PoLeCe}})$ at $\alpha=0.05$ is 587, which is much less than no-precision-corrected $\widehat{V}(\widehat{\pi}_{\text{EWM}})$ but substantially greater than $\widehat{V}(\widehat{\pi}_{\text{EWM}})-\widehat{q}_{0.95,\Pi}\widehat{s}(\widehat{\pi}_{\text{EWM}})$. Furthermore, the chosen optimal policies differ for females. EWM selects Health Pot for unmarried women and Safe Box for married women. In contrast, after accounting for standard errors, PoLeCe recommends Safe Box for unmarried women and Health Pot for married women.
@@ 3.2.3. Alcohol and Self-Control
Panel B. Selected Policy:
@@
Note. EWM refers to empirical welfare maximization, and PoLeCe corresponds to policy learning with confidence. In Panel A, the EWM point estimate and the PoLeCe with $\alpha=0.05$ are provided along the control group policy. In Panel B, selected treatment for each vector of (Baseline sober, Owns rickshaw) is given.
There are $|\mathcal{T}|=81=3^{4}$ possible treatment policies in total since there are three treatment levels and four possible values of the two covariates. Figure 3(b) plots the welfare estimates against their precision. Although the highest welfare appears with a relatively small standard error, this does not necessarily imply statistical significance. To implement PoLeCe at the level $\alpha=0.05$ and calculate $\widehat{q}_{0.95,\Pi}$, we apply the bootstrap procedure from Condition (B). Table 6 summarizes the empirical results. The EWM and PoLeCe estimates, along with their lower confidence bounds, are very similar and only slightly exceed those of the control group.
@@ 4. Conclusion
In this paper we have focused on the problem faced by a DM who wishes to choose from a menu of policies that which will deliver the highest welfare using imperfect sample estimates. In order to balance the estimated performance of each policy with its associated statistical uncertainty, we proposed and analyzed the properties of a class of risk-aware policies that make the precision/performance tradeoff explicit. Such policies solve the expected loss function of a risk-averse planner and achieve favorable regret properties. We proposed a specific rule from the class of risk-aware policies, namely the PoLeCe rule, which uses a data-dependent construction for balancing the inherent tradeoff between estimated performance and sample uncertainty, such that a lower confidence bound on the welfare obtained by the chosen policy is automatically provided. Application of the rule was demonstrated across four settings; one featuring the allocation of government funds across social programs using the MVPF, and three others featuring optimal treatment assignment. Intuitively, the approach offers a principled and robust way for a decision maker to hedge against selecting the best-performing policy in sample, when that policy exhibits large sample variation relative to other well-performing policies. Importantly, risk-aware policy rules allow policies with fractional allocations.
@@
A large body of work synthesized in Manski (2013, 2019, 2024) has advocated for greater acknowledgement and incorporation of uncertainty in planning problems. This paper contributes by proposing a principled way to acknowledge and incorporate statistical uncertainty into decision making. Statistical uncertainty is however only one of the many different types of uncertainty that may be present. [^20] For example, Manski (2019) discusses transitory uncertainty, permanent uncertainty, and conceptual uncertainty. The statistical uncertainty considered here is one type of permanent uncertainty. We have focused on settings in which the welfare of each policy is point identified, such that consistent estimates of the welfare and sample variance of the policies are available. Application of the concepts developed here to settings that feature so-called “deep uncertainty”, or ambiguity, would seem an important direction for further development given the large number of settings in which the mean performance of various policies is only credibly partially identified. [^21] There is now a long line of research on treatment choice when mean treatment performance is only partially identified, going back to Manski (2000). Recent research with references to the broader literature includes Russell (2020), Ishihara and Kitagawa (2021), Yata (2021), Christensen et al. (2023), Kido (2023), and Montiel Olea et al. (2023). This would require balancing statistical uncertainty with a collection of interval estimates for each policy’s performance.
@@
While the nuance required to formally develop such an approach is beyond the scope of the present paper, a rough prescription could be made based on an extension of the reporting guarantee established here. To see how, consider the definition of the PoLeCe rule in (7). The rule selects the policy that maximizes a $1-\alpha$ lower confidence band for the maximal achievable welfare. Construction of such a confidence band does not require that the optimal level of welfare or the optimal rule be point identified, and could be implemented using techniques developed in e.g. Chernozhukov et al. (2013) under partial identification. This is of course not the only way to balance statistical uncertainty and ambiguity. [^22] For instance, Ben-Michael et al. (2025) develops an alternative approach employing a minimax rule over a uniform two-sided confidence band for the conditional mean of potential outcomes in a setting with partial identification due to lack of overlap. A more thorough study of the performance of such an approach, and the implicit tradeoff between statistical uncertainty and ambiguity due to partial identification would seem a useful direction for future research.
@@ Appendix A Deferred Proofs
and by the standard calculation we have $\mathrm{E}\sup_{\pi\in{K}}Z_{\pi}\leq\sqrt{2\log|K|}$, where $|{K}|\leq p$ is the cardinality of ${K}$. ∎
@@
$\bar{\sigma}_{\Pi}:=\sqrt{n}\max_{\pi\in\Pi}\widehat{s}(\pi);\quad\widehat{\pi}:=\widehat{\pi}_{\text{RW}}(\widehat{k});$
@@
$V_{\max}-V(\widehat{\pi})$ $=$ $V_{\max}-L(\widehat{k})+L(\widehat{k})-V(\widehat{\pi})$
@@
Proof of Proposition 4
Claim (1). Let $v=\widehat{V}(\pi)$ and $s=\widehat{s}(\pi)$. Then
[…]
This maximization can be written in a Lagrangian form:
Proof of Proposition 5.
Observing that
and
the program becomes
@@ Appendix B Efficient Decision Frontier Computation
For the investment allocation problem, it may be infeasible to discretize $\Pi$ and apply the above algorithm to plot the efficiency frontier. Instead, we start with a fine grid for $k$ centered at $\hat{q}_{0.95,\Pi},$ and directly trace out the efficiency decision frontier by solving the corresponding optimization problem as in  (4). Appendix C provides computationally efficient way to solve such optimization and therefore to plot the efficiency frontier.
@@ Appendix C Gaussian Bootstrap for Investment Allocation Problem
$f(t)=\max_{\pi\in\Pi}\left(\pi^{\top}z-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}\right)$
when $Z^{*}=z$. The following lemma formalizes the properties of $f(t)$ necessary for this approach.
@@
Let $\Pi\subset\mathbb{R}^{d}$ be a convex, closed subset of the probability simplex, and let $\widehat{\Omega}$ be a positive-definite matrix. Define
@@
which is finite since $\Pi$ is bounded and closed. As $t\to\infty$, the term $-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}$ dominates for every $\pi\in\Pi$, leading to $f(t)\to-\infty$. Since $f(t)$ is continuous (as the pointwise maximum of continuous functions over a compact set), the Intermediate Value Theorem guarantees the existence of some $t^{*}\geq 0$ such that $f(t^{*})=0$.
@@ Appendix D Further Details on the MVPF Estimates
For ease of reference, the Policy Impacts Library (Hendren and Sprung-Keyser (2021)) descriptions of the social programs corresponding to each MVPF estimate in Tables 2 and 3 are provided below.